OpenStudy (anonymous):

If tan x = .76 and cos x = .22, what is sin x?

OpenStudy (ash2326):

We know that \[\tan x= \frac{\sin x}{\cos x}\] so \[\sin x=\tan x \times \cos x\] Just plugin the values and voila!! I think you can do now, can't you?

OpenStudy (anonymous):

i got 0.1672 is that right ?

OpenStudy (ash2326):

yeah, you're right, Good work

OpenStudy (anonymous):

thankks^.^